OpenStudy (anonymous):

((3)^-2)^x=√(3) Find x. Also, if possible, please explain how you got the answer. Thanks!

OpenStudy (anonymous):

((3)^-2)^x=√(3) taking log we get log((3)^(-2x)=log√(3) -2xlog((3)=(1/2)log(3) -2x=(1/2)(log3)/(log3) -2x=1/2 x=-1/4 ans